# Fast Transients from Magnetic Disks Around Non-Spinning Collapsar Black Holes

### v2

@@ Abstract
Most black holes (BHs) formed in collapsing stars have low spin, though some are expected to acquire a magnetic accretion disk during the collapse. While such BH disks can launch magnetically driven winds, their physics and observational signatures have remained unexplored. We present global 3D general relativistic magnetohydrodynamic simulations of collapsing stars that form slowly spinning BHs with accretion disks. As the disk transitions to a magnetically arrested state, it drives mildly relativistic, wobbling, collimated magnetic outflows through two mechanisms: steady outflows along vertical magnetic field lines (“Blandford-Payne jets”) and magnetic flux eruptions. With an isotropic-equivalent energy of $E_{\rm iso}\approx 10^{52}\,{\rm erg}$, exceeding that of relativistic jets from BHs with spin $a\lesssim 0.25$, the disk outflows unbind the star, ultimately capping the final BH mass at $M_{\rm BH}\approx 4\,M_{\odot}$. Once the outflows emerge from the star, they produce mildly relativistic shock breakout, cooling, and ${}^{56}{\rm Ni}$-decay emission. Our cooling emission estimates suggest a bright near-ultraviolet and optical signal at absolute magnitude $M_{\rm AB}\approx-16$ lasting for several days. This indicates that disk winds could be responsible for the first peak in the double-peaked light curves observed in Type Ib/c supernovae (SNe) or power another class of transients. The detection rate in the upcoming Rubin Observatory and ULTRASAT/UVEX will enable us to differentiate between competing models for the origin of the first SN peak and provide constraints on the physics and formation rate of accretion disks in core-collapse SNe.
@@ 2 Numerical setup
We perform a series of 3D GRMHD collapsar simulations using the GPU-accelerated code h-amr (Liska et al., 2022), employing an ideal equation of state featuring an adiabatic index of $\gamma=4/3$. The simulations follow the setup detailed in Gottlieb et al. (2022a), featuring a BH of mass $\,{M_{\rm BH}}=4\,\,{\rm M_{\odot}}$ embedded within a Wolf-Rayet star with a stellar radius of $R_{*}=4\times 10^{10}\,\,{\rm cm}$ and mass $M_{*}\approx 14\,\,{\rm M_{\odot}}$. We vary the dimensionless spin parameter, $a$, as outlined in Table 1. The only model in which the BH spin might appreciably change during the simulation is $\alpha 0\omega\sigma HR$, which runs for $23.3\,\,{\rm s}$ with an initial BH spin of $a=0$. However, we do not evolve the BH spin with accretion, as the outflow physics remains similar between BHs with spins of $a=0$ and $a_{\rm eq}=0.1$, as we will demonstrate.
@@
Model name $a$ Disk MAD $t_{f}\,[\,{\rm s}]$
$a0$ 0.0 No No 3.5
$a0\omega$ 0.0 Yes No 1.5
$a0\sigma$ 0.0 No Yes 0.7
$a0\omega\sigma$ 0.0 Yes Yes 1.5
$a0\omega\sigma HR$ 0.0 Yes Yes 23.3
$a1\omega\sigma$ 0.1 Yes Yes 1.4
$a2\omega\sigma$ 0.2 Yes Yes 1.5
$a3\omega\sigma$ 0.3 Yes Yes 1.5
$a5\omega\sigma$ 0.5 Yes Yes 0.6
$a9$ 0.9 No No 2.2
$a9\omega$ 0.9 Yes No 1.2
$a9\sigma$ 0.9 No Yes 0.9
$a9\omega\sigma$ 0.9 Yes Yes 1.2
Table 1: Model parameters. The model name structure is as follows: $a\#$ represents the dimensionless BH spin parameter, where #$\#=10\times a$; $\omega$ is included (omitted) for models that (do not) form disks with $\omega_{0}=50\,\,{\rm s}^{-1}$ ($\omega_{0}=0.5\,\,{\rm s}^{-1}$); $\sigma$ is shown when magnetic fields are included; $HR$ denotes a high-resolution run. $t_{f}$ is the final time of the simulation.
@@ 3.1.3 Disks without a strong magnetic field: Accretion shock
In scenarios lacking magnetic fields – likely when the disk is not initially present to preserve the PNS’s magnetic field (Gottlieb et al., 2024b) – the disk transports angular momentum outward through hydrodynamic viscous forces. This process energizes an expanding accretion shock (see models $a0\omega$, $a9\omega$). While such shocks are also present in magnetically driven outflows, their contribution is more prominent when magnetic forces are absent. The shock exhibits an $m=1$ mode, similar to accretion shocks observed in core-collapse SN (CCSN) simulations (Blondin et al., 2003; Blondin & Mezzacappa, 2006).
@@ 3.2 Outflow comparison
Figure 2 shows the time evolution of various quantities at the BH horizon in simulations that form accretion disks. Panel (a) compares the outflow launching efficiency, defined as
@@
Panel (b) shows the mass accretion rate, $\dot{M}$, onto the BH across different models, demonstrating that the strong outflows, produced by higher spin BHs, slightly suppress accretion. $a0\omega\sigma HR$ exhibits a sharper decrease in accretion at $t\gtrsim 0.5\,\,{\rm s}$ due to this feedback mechanism. Panel (c) depicts the radial power at the horizon [Eq. (7)], indicating total outflow energies of $E\sim 10^{51}-10^{52}\,\,{\rm erg}$.
Figure 3 compares the empirical expression for the spin efficiency of BZ jets in Eq. (5) (blue curve) with the time-averaged launching efficiency of various outflows in our simulations (dots). For each spin, the maximum power is achieved in simulations with both strong magnetic fields and fast rotation (black and khaki dots). The empirical equation is consistent with the numerical values of simulations with $a\gtrsim 0.3$, implying that the BZ mechanism dominates the outflows. Simulations of magnetic disks with $a\lesssim 0.3$ exhibit outflows with considerably higher efficiency than the prediction by the blue curve. This indicates that an additional ejection mechanism is at play, driven by the accretion disk. For example, when $a=0$, the efficiency $\eta\approx 3\%$ is equivalent to BZ outflows from a BH with spin $a_{\rm crit}\approx 0.25$ (corresponds to $\eta(a)\approx 3\%$). This suggests that outflows from BHs with $a\lesssim 0.25$ are dominated by non-BZ outflows, consistent with McKinney et al. (2012); Narayan et al. (2012); Tchekhovskoy & McKinney (2012). Simulations without rotation (turquoise) prevent disk formation, leading to brief, jittering jet launches due to the accumulation of flux on the BH. When magnetic flux is absent (red), the resulting hydrodynamic outflows are weaker by orders of magnitude compared to the magnetically driven outflows.
Figure 3: The time-averaged outflow efficiency in simulations with strong magnetic fields and rapid rotation (black dots) follows Eq. (5) (blue curve) for $a\gtrsim 0.3$. For $a\lesssim 0.3$, the outflow efficiency is dominated by disk outflows rather than the BH. BHs with weak magnetic fields (red) or without accretion disks (turquoise) exhibit a low launching efficiency.
@@ 4.1 Explosion physics
Fig. 4 shows meridional (top row) and equatorial (bottom row) maps of the region near the BH during a magnetic flux eruption in the non-spinning BH model $a0\omega\sigma HR$. Magnetic pressure accumulates in low-density regions near the BH [panels (b) and (d)] and is transported radially outward. Ultimately, the plasma is expelled from the disk [see velocity streamlines in panel (a)] along the vertical field lines [white lines in panel (c)]. In addition to the collimated flux eruptions, magnetically driven winds are constantly launched along the vertical field lines as BP outflows (see e.g., Musolino et al., 2024). However, during the flux eruptions, most of the vertical power of the outflow,
@@
We find that the magnetic flux eruptions drive mildly relativistic ($\Gamma\approx 2$) outflows that occur every $\sim 50\,\,{\rm ms}$ for a duration of $\sim 10\,\,{\rm ms}$, consistent with the reconnection rate in collisional plasma (Bhattacharjee et al., 2009; Uzdensky et al., 2010). This indicates that flux eruptions are present for $\sim 20\%$ of the time. Fig. 4(e),(f) demonstrates that during an eruption, the vertical power from the flux eruption constitutes almost all the disk power. This indicates that the integrated energy from these flux eruptions may dominate over the BP mechanism, making flux eruptions the primary driver of the disk outflows.
Figure 5 presents a 3D rendering of the magnetization of the ejecta after breakout from the star (white dotted sphere). The mildly collimated, massive disk outflows (red) at $v\approx 0.15c$ are embedded within a quasi-spherical, shocked, double-lobed structure (blue) moving at $v\lesssim 0.4c$. As the disk-driven outflows interact with the infalling stellar material, they form a double-shocked cocoon layer. The shocked cocoon material falls back onto the disk, tilting it away from the axis of rotation and driving the wobbling motion (see Gottlieb et al., 2022b). The top panel of Figure 6 depicts the angular distribution of the isotropic equivalent energy of the outflows outside the star. It illustrates that the $E\approx 10^{51.5}\,\,{\rm erg}$ ejecta is concentrated within a characteristic opening angle of $\theta_{e}\approx\pi/4$, resulting in isotropic equivalent energy of $E_{\rm iso}\approx 10^{52}\,\,{\rm erg}$.
@@ 4.2 Electromagnetic signatures
ULTRASAT will observe $\sim 6800$ deg2 of the sky during its low-cadence survey mode, corresponding to $f_{\Omega}\approx 1/6$ of the sky. If all Type Ic-BL SNe harbor such accretion disks, and assuming emission only within $\theta_{e}$, then the ULTRASAT detection rate is $\sim(2\pi/3)\theta_{e}^{2}D^{3}\mathcal{R}_{\rm Ic-BL}f_{\Omega}\sim 30\,{\rm yr}^{-1}$. A similar calculation for Rubin Observatory yields a detection rate of $\gtrsim 10^{4}\,{\rm yr}^{-1}$. However, these rates strongly depend on the disk properties such as magnetization, and the uncertain rate of disk formation in Type Ic-BL SNe and other CCSN classes.
5 Conclusions
@@
The disk outflows are driven by a comparable contribution of vertical magnetic field lines threading the accretion disk (BP jets) and magnetic flux eruptions from regions of high magnetic pressure near the BH. Feedback from the cocoon, which is generated by the disk winds, induces a disk tilt that causes the outflows to wobble. With an isotropic-equivalent energy of $E_{\rm iso}\approx 10^{52}\,\,{\rm erg}$, these outflows explode the star, leaving merely $\sim 1\,\,{\rm M_{\odot}}$ for BH mass growth. After breakout from the star, the disk outflows will generate multiple types of emission across the spectrum: shock breakout, cooling envelope, ${}^{56}{\rm Ni}$ decay, and synchrotron emission. We find that their cooling emission will be detectable out to hundreds of Mpc by ULTRASAT and $D>{\rm Gpc}$ by UVEX and the Rubin Observatory. The detection rate of these events will provide valuable constraints on accretion disk physics and its formation rate in CCSNe.
@@
We thank Amir Levinson for useful discussions. J.B. acknowledges the support from the Simons Foundation Presidents Discretionary Fund. O.G. is supported by the Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085).

### v3

@@ Abstract
Most black holes (BHs) formed in collapsing stars have low spin, though some are expected to acquire a magnetic accretion disk during the collapse. While such BH disks can launch magnetically driven winds, their physics and observational signatures have remained unexplored. We present global 3D general relativistic magnetohydrodynamic simulations of collapsing stars that form slowly spinning BHs with accretion disks. As the disk transitions to a magnetically arrested state, it drives mildly relativistic, wobbling, collimated magnetic outflows through two mechanisms: steady outflows along vertical magnetic field lines (“Blandford-Payne jets”) and magnetic flux eruptions. With isotropic-equivalent energy of $E_{\rm iso}\approx 10^{52}\,{\rm erg}$, exceeding that of relativistic jets from BHs with spin $a\lesssim 0.25$, the disk outflows unbind the star, ultimately capping the final BH mass at $M_{\rm BH}\approx 4\,M_{\odot}$. Once the outflows emerge from the star, they produce mildly relativistic shock breakout and cooling emission. Our cooling emission estimates suggest a bright near-ultraviolet and optical signal at absolute magnitude $M_{\rm AB}\approx-16$ lasting for several days. This indicates that disk winds could be responsible for the first peak in the double-peaked light curves observed in Type Ib/c supernovae (SNe) or power another class of transients. The detection rate in the upcoming Rubin Observatory and ULTRASAT/UVEX will enable us to differentiate between competing models for the origin of the first SN peak and provide constraints on the physics and formation rate of accretion disks in core-collapse SNe.
@@ 2 Numerical setup
We perform a series of 3D GRMHD collapsar simulations using the GPU-accelerated code h-amr (Liska et al., 2022), employing an ideal equation of state featuring an adiabatic index of $\gamma=4/3$. The simulations follow the setup detailed in Gottlieb et al. (2022a), featuring a BH of mass $\,{M_{\rm BH}}=4\,\,{\rm M_{\odot}}$ embedded within a Wolf-Rayet star with a stellar radius of $R_{*}=4\times 10^{10}\,\,{\rm cm}$ and mass $M_{*}\approx 14\,\,{\rm M_{\odot}}$. We vary the dimensionless spin parameter, $a$, as outlined in Table 1. The only model in which the BH spin might appreciably change during the simulation is $\alpha 0\omega\sigma HR$, which runs for $23.3\,\,{\rm s}$ with an initial BH spin of $a=0$. However, we do not evolve the BH spin with accretion, as the outflow physics remains similar between BHs with spins of $a=0$ and $a_{\rm eq}=0.1$, as we will demonstrate. Due to computational time limitations, all other models were run until their outflow type (see below) could be determined.
@@
Model name $a$ Disk MAD Outflow $t_{f}\,[\,{\rm s}]$
$a0$ 0.0 No No None 3.5
$a0\omega$ 0.0 Yes No AS 1.5
$a0\sigma$ 0.0 No Yes None 0.7
$a0\omega\sigma$ 0.0 Yes Yes APO 1.5
$a0\omega\sigma HR$ 0.0 Yes Yes APO 23.3
$a1\omega\sigma$ 0.1 Yes Yes APO 1.4
$a2\omega\sigma$ 0.2 Yes Yes APO 1.5
$a3\omega\sigma$ 0.3 Yes Yes BZ 1.5
$a5\omega\sigma$ 0.5 Yes Yes BZ 0.6
$a9$ 0.9 No No None 2.2
$a9\omega$ 0.9 Yes No AS 1.2
$a9\sigma$ 0.9 No Yes JJ 0.9
$a9\omega\sigma$ 0.9 Yes Yes BZ 1.2
Table 1: Model parameters. The model name structure is as follows: $a\#$ represents the dimensionless BH spin parameter, where #$\#=10\times a$; $\omega$ is included (omitted) for models that (do not) form disks with $\omega_{0}=50\,\,{\rm s}^{-1}$ ($\omega_{0}=0.5\,\,{\rm s}^{-1}$); $\sigma$ is shown when magnetic fields are included; $HR$ denotes a high-resolution run. The outflow types are accretion shock (AS), accretion-powered outflow (APO), BZ jets (BZ), and jittering jets (JJ). $t_{f}$ is the final time of the simulation.
@@ 3.1.1 No disk: Jittering jets
Over time, the propagation of outflows within a collapsing star can locally enhance angular momentum in certain regions. When these regions reach the BH, they may form an accretion disk. However, this accretion disk will subsequently ingest gas with opposite angular momentum, which destabilizes the disk and the outflows, leading to the formation of a rocking accretion disk (RAD; Lalakos2024). The RAD formation may gradually align the disk and BH spin, potentially stabilizing the jet and producing more coherent, sustained outflows.
@@ 3.1.2 Magnetized Kerr BHs with disks: BZ-jets (lGRBs)
This highlights the critical role of BH spin in determining the efficiency and power of BZ jet launching in collapsars, and shaping the observable characteristics of lGRBs.
@@ 3.1.3 Disks without a strong magnetic field: Accretion shock
In scenarios lacking magnetic fields – likely when the disk is not initially present to preserve the PNS’s magnetic field (Gottlieb et al., 2024b) – the disk transports angular momentum outward through hydrodynamic viscous forces. This process energizes an expanding accretion shock (see models $a0\omega$, $a9\omega$). While such shocks are also present in magnetically driven outflows, their contribution is more prominent when magnetic forces are absent. The shock exhibits an $m=1$ mode, similar to accretion shocks observed in core-collapse SN (CCSNe) simulations (Blondin et al., 2003; Blondin & Mezzacappa, 2006).
@@ 3.2 Outflow comparison
Figure 2 shows the time evolution of various quantities at the BH horizon in simulations that form accretion disks. Panel (a) of Figure 2 compares the outflow launching efficiency, defined as
@@
Panel (b) of Figure 2 shows the mass accretion rate, $\dot{M}$, onto the BH across different models, demonstrating that the strong outflows, produced by higher spin BHs, slightly suppress accretion. $a0\omega\sigma HR$ exhibits a sharper decrease in accretion at $t\gtrsim 0.5\,\,{\rm s}$ due to this feedback mechanism. Panel (c) of Figure 2 depicts the radial power at the horizon [Eq. (7)], indicating total outflow energies of $E\sim 10^{51}-10^{52}\,\,{\rm erg}$.
Figure 3 compares the empirical expression for the spin efficiency of BZ jets in Eq. (5) (blue curve) with the time-averaged launching efficiency of various outflows in our simulations. For each spin, the maximum power is achieved in simulations with both strong magnetic fields and fast rotation (red dots and black X). The empirical equation is consistent with the numerical values of simulations with $a\gtrsim 0.3$, implying that the BZ mechanism dominates the outflows. Simulations of magnetic disks with $a\lesssim 0.3$ exhibit outflows with considerably higher efficiency than the prediction by the blue curve. This indicates that an additional ejection mechanism is at play, driven by the accretion disk. For example, when $a=0$, the efficiency $\eta\approx 3\%$ is equivalent to BZ outflows from a BH with spin $a_{\rm crit}\approx 0.25$ (corresponds to $\eta(a)\approx 3\%$). This suggests that outflows from BHs with $a\lesssim 0.25$ are dominated by non-BZ outflows, consistent with McKinney et al. (2012); Narayan et al. (2012); Tchekhovskoy & McKinney (2012). Simulations without rotation (turquoise diamonds) prevent disk formation, leading to brief, jittering jet launches due to the accumulation of flux on the BH. When magnetic flux is absent (green square), the resulting hydrodynamic outflows are weaker by orders of magnitude compared to the magnetically driven outflows.
Figure 3: The time-averaged outflow efficiency in simulations with strong magnetic fields and rapid rotation (red dots) follows Eq. (5) (blue curve) for $a\gtrsim 0.3$. For $a\lesssim 0.3$, the outflow efficiency is dominated by disk outflows rather than the BH. BHs with weak magnetic fields (green square) or without accretion disks (turquoise diamonds) exhibit a low launching efficiency.
@@ 4.1 Explosion physics
Fig. 4 shows meridional (top row) and equatorial (bottom row) maps of the region near the BH during a magnetic flux eruption in the non-spinning BH model $a0\omega\sigma HR$. Magnetic pressure accumulates in low-density regions near the BH [panels (b) and (d)] and is transported radially outward. Ultimately, the plasma is expelled from the disk [see velocity streamlines in panel (a)] along the vertical field lines [white lines in panel (c)]. In addition to the collimated flux eruptions, magnetically driven winds are constantly launched along thepoloidal field lines as BP outflows (see e.g., Musolino et al., 2024). However, during the flux eruptions, most of the vertical power of the outflow,
@@
Fig. 4(e),(f) demonstrates that during an eruption, the vertical power from the flux eruption constitutes almost all the disk power. This indicates that the integrated energy from these flux eruptions may dominate over the BP mechanism, making flux eruptions the primary driver of the disk outflows.
Figure 5 presents a 3D rendering of the magnetization of the ejecta after breakout from the star (white dotted sphere). The mildly collimated, massive disk outflows (red) at $v\approx 0.15c$ are embedded within a quasi-spherical, shocked, double-lobed structure (blue) moving at $v\lesssim 0.4c$. As the disk-driven outflows interact with the infalling stellar material, they form a double-shocked cocoon layer. The shocked cocoon material falls back onto the disk, tilting it away from the axis of rotation and driving the wobbling motion. Our temporal resolution has prevented us from conducting a detailed analysis of the wobbling origin (see Gottlieb et al., 2022b, for analysis and discussion). The top panel of Figure 6 depicts the angular distribution of the isotropic equivalent energy of the outflows outside the star. It illustrates that the $E\approx 10^{51.5}\,\,{\rm erg}$ ejecta is concentrated within a characteristic opening angle of $\theta_{e}\approx\pi/4$, resulting in isotropic equivalent energy of $E_{\rm iso}\approx 10^{52}\,\,{\rm erg}$.
@@ 4.2 Electromagnetic signatures
ULTRASAT will observe $\sim 6800$ deg2 of the sky during its low-cadence survey mode, corresponding to $f_{\Omega}\approx 1/6$ of the sky. If all Type Ic-BL SNe harbor such accretion disks, and assuming emission only within $\theta_{e}$, then the ULTRASAT detection rate is $\sim(2\pi/3)\theta_{e}^{2}D^{3}\mathcal{R}_{\rm Ic-BL}f_{\Omega}\sim 30\,{\rm yr}^{-1}$. A similar calculation for Rubin Observatory yields a detection rate of $\gtrsim 10^{4}\,{\rm yr}^{-1}$. However, these rates strongly depend on the disk properties such as magnetization, and the uncertain rate of disk formation in Type Ic-BL SNe and other CCSNe classes.
5 Summary and Discussion
@@
The disk outflows are driven by a comparable contribution of poloidal magnetic field lines threading the accretion disk (BP jets) and magnetic flux eruptions from regions of high magnetic pressure near the BH. Feedback from the cocoon, which is generated by the disk winds, induces a disk tilt that causes the outflows to wobble. With an isotropic-equivalent energy of $E_{\rm iso}\approx 10^{52}\,\,{\rm erg}$, these outflows explode the star, leaving merely $\sim 1\,\,{\rm M_{\odot}}$ for BH mass growth. After breakout from the star, the disk outflows will generate multiple types of emission across the spectrum: shock breakout, cooling envelope, ${}^{56}{\rm Ni}$ decay, and synchrotron emission. We find that their cooling emission will be detectable out to hundreds of Mpc by ULTRASAT and $D>{\rm Gpc}$ by UVEX and the Rubin Observatory. The detection rate of these events will provide valuable constraints on accretion disk physics and its formation rate in CCSNe.
@@
We thank Amir Levinson and Aris Lalakos for useful comments. J.B. acknowledges the support from the Simons Foundation Presidents Discretionary Fund. O.G. is supported by the Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085).